# Jaipur-302017, INDIA

### v1

@@ 1 Introduction
The study of spectral and critical properties of random matrix models has become a central interdisciplinary topic, as these models naturally arise in quantum field theory, condensed matter physics, number theory, and statistics Mehta1991; log-gases_2010; Baik2016; guhr_random-matrix_1998. Unitary matrix models appear in a wide variety of physical systems, such as gauge theories gross_possible_1980; aharony2004; dumitru2005; hands_qcd_2010, two-dimensional Quantum Gravity brezin_exactly_1990; douglas_strings_1990; gross_nonperturbative_1990, the Ising model and disordered mesoscopic conductors mccoy_two-dimensional_1973; muttalib_new_1993; deift2013, as well as topological field and string theories vleeshouwers_topological_2021. They are also connected to the Painlevé equations and their $\tau$-functions forrester2001; szabo_two-dimensional_2012. Simpler/limiting versions of unitary matrix models have been investigated for decades by both mathematicians and physicists; see Refs. mandal_phase_1990; russo_phases_2020; santilli_exact_2020; santilli_multiple_2022 and the references therein. We also refer to the excellent literature reviews Refs. rossi_large-n_1998 and marino_houches_2005.
Depending on the system, the matrix model description may be non-perturbative, which has its own obvious advantages. In fact, in certain cases the non-perturbative structure can be determined in complete detail ahmed_transmutation_2017; Anees2018, and in such cases matrix models serve as a wonderful test lab to explore resurgent asymptotics. Additionally, matrix models have attracted attention in recent years due to their relevance to systems that exhibit the sign problem, which obstructs the use of standard techniques such as Monte-Carlo in lattice gauge theory hands_simulating_2007. Recent progress in resolving the sign problem has involved methods such as complex Langevin dynamics aarts2010; boguslavski2025; klauder_coherent-state_1984 and the Lefschetz-thimble approach fujii_monte_2015; cristoforetti2013; nishimura2024.
@@
This potential is complex valued. We note here that a large number of unitary models studied in the literature have a real potential. As the potential (3) is complex and the model (1) is an SU($N$) model, the relation $\langle U\rangle=\langle U^{-1}\rangle$ may not necessarily hold for our model, whereas in the case of U($N$) models with real potentials it always holds. The branch points of the logarithm in (3) are Fisher-Hartwig singularities (see Ref. fahs_uniform_2021 and the references therein), and their existence implies a rich phase structure.
The $\gamma\to\infty$ limit of this model has been studied in Ref. hands_qcd_2010, where it describes the statistics of a two-level QCD (Quantum Chromodynamics) system restricted to a small sphere. In the triple-scaling limit $\sigma\to 0,\alpha\to 0,\gamma\to\infty$ while keeping $\alpha\sigma$ and $\sigma/\gamma$ equal and constant, the model reduces to the SU($N$) version of the well-known Gross-Witten-Wadia model gross_possible_1980. Thus our model can be viewed as a generalization of several existing models.
@@ 2 Formalism
The problem now is to find the spectral function $\rho(z)$ and its support $\mathcal{C}$ given the potential (3). This is a Reimann-Hilbert problem brezin1978; marino_houches_2005, and it can be solved in a rather straight-forward manner if we restrict to solutions in which $\mathcal{C}$ is a single continuous curve. Such a solution is called a single cut solution. If the potential is real and model is a U($N$) matrix model (i.e. $\mathcal{N}=0$), then the saddle points are found exclusively on the unit circle and the support $\mathcal{C}$ is simply the unit circle or an arc of a unit circle. In the present case the saddle points may be found off the unit circle.
@@ 3 Ungapped phases
If our model results from a lattice gauge theory, then these moments are related to the winding Wilson loops over a single plaquette okuyama2017. We observe that
@@
which indicates the presence of the sign problem. This behavior is completely expected, as our model’s potential is complex, in addition to having an SU($N$) geometry. In the Gross-Witten-Wadia model gross_possible_1980, the single-winding Wilson loop ($W_{1}$) is related the string tension $\tau$ and the lattice spacing $a$ as $W_{1}=e^{-a^{2}\tau}$. If we assume that this continues to hold if our model is connected to a lattice gauge theory, then the beta function has the expression
$\beta=-2W_{1}\dfrac{\ln W_{1}}{\partial_{\sigma}W_{1}}.$ (25)
For phase A this yields $\beta=-2\sigma\ln(\sigma/\gamma)$.
@@
Beta function ${-2\sigma}\ln{({\sigma}/{\gamma})}$ ${-2\sigma}\ln{\left(\dfrac{\sigma}{\alpha}+\dfrac{\sigma}{\gamma}\right)}$
@@
Beta function ${-2\sigma}\ln{({\sigma}/{\alpha})}$ Undefined
@@ 4 Gapped Phase
By using the Sokhotski-Plemelj formula muskhelishvili2013, we find the two main expressions which make the resolvent a useful quantity. The first relates the potential to the resolvent:
@@
This result is similar to a theorem by Tricomi tricomi1985.
@@
By deforming the contour $\mathcal{C}^{*}$ as shown in Fig. 3, we find the SU($N$) contraint to be
@@ 5 Phase transitions
Lastly, as mentioned in the introduction, our model converges to the low temperature two-level QCD model of Ref. hands_qcd_2010 if we relate the parameters $\alpha,\gamma$ to the temperature $T$ and the chemical potential $\mu$ as follows
@@
where $\varepsilon$ is the energy difference between the two levels. In this case, $\sigma$ is proportional to the number of quark flavors, $\mathscr{N}$ becomes the fermion number, and the ungapped and gapped phases have the interpretation of confined and deconfined phases. In the exact zero temperature limit $\gamma$ approaches infinity, and one of the log terms in the potential (3) can be dropped, which is exactly what the authors of hands_qcd_2010 have done. Keeping both terms produces the phase diagram shown in Fig. 8, which differs from that of hands_qcd_2010 only at large temperatures. Only two confined phases exist in this diagram, as the other two correspond to negative temperature. We also note that as $\sigma$ increases, the deconfined phase expands, while the confined phase shrinks to smaller values of $T$. The three phases exist at every finite value of $\sigma$. Lastly, it appears that a direct confined to confined transition is allowed at $T=0$. This apparent contradiction with our earlier discussion is explained by the fact that the potential (3) does not really cover the $T=0$ case as the parameters $\alpha$ and $\gamma$ are ill-defined at $T=0$.

### v2

@@ 1 Introduction
The study of spectral and critical properties of random matrix models has become a central interdisciplinary topic, as these models naturally arise in quantum field theory, condensed matter physics, number theory, and statistics 27; 23; 5; T. Guhr, A. Müller–Groeling, and H. A. Weidenmüller (1998). Unitary matrix models appear in a wide variety of physical systems, such as gauge theories Gross and Witten (1980); Aharony et al. (2004); Dumitru et al. (2005); Hands et al. (2010), two-dimensional Quantum Gravity Brézin and Kazakov (1990); Douglas and Shenker (1990); Gross and Migdal (1990), the Ising model and disordered mesoscopic conductors McCoy and Wu (1973); Muttalib et al. (1993); Deift et al. (2013), as well as topological field and string theories Vleeshouwers and Gritsev (2021). They are also connected to the Painlevé equations and their $\tau$-functions J. Forrester and S. Tierz (2001); Szabo and Tierz (2012). Simpler/limiting versions of unitary matrix models have been investigated for decades by both mathematicians and physicists; see Refs. Mandal (1990); Russo (2020); Santilli and Tierz (2020, 2022) and the references therein. We also refer to the excellent literature reviews Refs. Rossi et al. (1998) and Marino (2005).
Depending on the system, the matrix model description may be non-perturbative, which has its own obvious advantages. In fact, in certain cases the non-perturbative structure can be determined in complete detail Ahmed and Dunne (2017, 2018), and in such cases matrix models serve as a wonderful test lab to explore resurgent asymptotics. Additionally, matrix models have attracted attention in recent years due to their relevance to systems that exhibit the sign problem, which obstructs the use of standard techniques such as Monte-Carlo in lattice gauge theory Hands (2007). Recent progress in resolving the sign problem has involved methods such as complex Langevin dynamics Aarts et al. (2010); Boguslavski et al. (2025); Klauder (1984) and the Lefschetz-thimble approach Fujii et al. (2015); Cristoforetti et al. (2013); Nishimura et al. (2024).
@@
This potential is complex valued. We note here that a large number of unitary models studied in the literature have a real potential. As the potential (3) is complex and the model (1) is an SU($N$) model, the relation $\langle U\rangle=\langle U^{-1}\rangle$ may not necessarily hold for our model, whereas in the case of U($N$) models with real potentials it always holds. The branch points of the logarithm in (3) are Fisher-Hartwig singularities (see Ref. Fahs (2021) and the references therein), and their existence implies a rich phase structure.
The $\gamma\to\infty$ limit of this model has been studied in Ref. Hands et al. (2010), where it describes the statistics of a two-level QCD (Quantum Chromodynamics) system restricted to a small sphere. In the triple-scaling limit $\sigma\to 0,\alpha\to 0,\gamma\to\infty$ while keeping $\alpha\sigma$ and $\sigma/\gamma$ equal and constant, the model reduces to the SU($N$) version of the well-known Gross-Witten-Wadia model Gross and Witten (1980). Thus our model can be viewed as a generalization of several existing models.
@@ 2 Formalism
The problem now is to find the spectral function $\rho(z)$ and its support $\mathcal{C}$ given the potential (3). This is a Reimann-Hilbert problem Brézin et al. (1978); Marino (2005), and it can be solved in a rather straight-forward manner if we restrict to solutions in which $\mathcal{C}$ is a single continuous curve. Such a solution is called a single cut solution. If the potential is real and model is a U($N$) matrix model (i.e. $\mathcal{N}=0$), then the saddle points are found exclusively on the unit circle and the support $\mathcal{C}$ is simply the unit circle or an arc of a unit circle. In the present case the saddle points may be found off the unit circle.
@@ 3 Ungapped phases
If our model results from a lattice gauge theory, then these moments are related to the winding Wilson loops over a single plaquette Okuyama (2017). We observe that
@@
which indicates the presence of the sign problem. This behavior is completely expected, as our model’s potential is complex, in addition to having an SU($N$) geometry. In the Gross-Witten-Wadia model Gross and Witten (1980), the single-winding Wilson loop ($W_{1}$) is related the string tension $\tau$ and the lattice spacing $a$ as $W_{1}=e^{-a^{2}\tau}$. If we assume that this continues to hold if our model is connected to a lattice gauge theory, then the beta function of $1/\sigma$ has the expression
$\beta=\dfrac{2W_{1}}{\sigma^{2}}\dfrac{\ln W_{1}}{\partial_{\sigma}W_{1}}.$ (25)
For phase A this yields $\beta=\dfrac{2}{\sigma}\ln{\dfrac{\sigma}{\gamma}}$.
@@
Beta function $\dfrac{2}{\sigma}\ln{\dfrac{\sigma}{\gamma}}$ $\dfrac{2}{\sigma}\ln{\left(\dfrac{\sigma}{\alpha}+\dfrac{\sigma}{\gamma}\right)}$
@@
Beta function $\dfrac{2}{\sigma}\ln{\dfrac{\sigma}{\alpha}}$ Undefined
@@ 4 Gapped Phase
By using the Sokhotski-Plemelj formula Muskhelishvili (2013), we find the two main expressions which make the resolvent a useful quantity. The first relates the potential to the resolvent:
@@
This result is similar to a theorem by Tricomi G. Tricomi (1985).
@@
By deforming the contour $\mathcal{C}^{*}$ as shown in Fig. 3, we find the SU($N$) constraint to be
@@ 5 Phase transitions
Lastly, as mentioned in the introduction, our model converges to the low temperature two-level QCD model of Ref. Hands et al. (2010) if we relate the parameters $\alpha,\gamma$ to the temperature $T$ and the chemical potential $\mu$ as follows
@@
where $\varepsilon$ is the energy difference between the two levels. In this case, $\sigma$ is proportional to the number of quark flavors, $\mathscr{N}$ becomes the fermion number, and the ungapped and gapped phases have the interpretation of confined and deconfined phases. In the exact zero temperature limit $\gamma$ approaches infinity, and one of the log terms in the potential (3) can be dropped, which is exactly what the authors of Hands et al. (2010) have done. Keeping both terms produces the phase diagram shown in Fig. 8, which differs from that of Hands et al. (2010) only at large temperatures. Only two confined phases exist in this diagram, as the other two correspond to negative temperature. We also note that as $\sigma$ increases, the deconfined phase expands, while the confined phase shrinks to smaller values of $T$. The three phases exist at every finite value of $\sigma$. Lastly, it appears that a direct confined to confined transition is allowed at $T=0$. This apparent contradiction with our earlier discussion is explained by the fact that the potential (3) does not really cover the $T=0$ case as the parameters $\alpha$ and $\gamma$ are ill-defined at $T=0$.